# A review of Bayesian sensor-based estimation and uncertainty quantification of aerodynamic flows

### v2

@@ 2.3.1 Sampling the posterior
So how do we proceed? If our goal is just to find the expected state given the true measurement, then we could note that this is the $\mathsf{x}$ that minimizes (22). This observation effectively frames this as an optimization problem, called the maximum a posterior (MAP) estimate, and we could use optimization tools to solve it. However, our goal is also to characterize the posterior uncertainty, so we still seek the posterior distribution in some form. We have a few options. If our prior distribution is already reasonably precise, then we do not have to move very far from it to obtain the posterior. In such a case we can linearize the observation operator about the prior mean and then use the Kalman update that we presented in the previous section, taking $H\approx\nabla h(\mu_{\mathsf{x}})$. We will revisit this idea later when we discuss sequential estimation. However, if our prior has a lot of uncertainty, or we suspect that the solution $\mathsf{x}$ to (1) is not unique, then we need a more thorough strategy, and we can get this by sampling the posterior. Sampling from a distribution provides us with an indirect tool for characterizing it: with an ensemble $\{\mathsf{x}^{i}\}_{i=1}^{M}$ of $M$ independent samples, each drawn from the posterior distribution $\mathsf{x}^{i}\sim\pi(\mathsf{x}|\mathsf{y}^{\ast})$, we can approximately compute expected values by summing over the samples,
@@
We can also use the samples to develop approximate probabilistic models of the posterior, $\tilde{\pi}(\mathsf{x}|\mathsf{y})$, to do further work, e.g., to serve as a generative model for drawing new instances of the state. There are many approaches for carrying out this approximation task, most of which consist of minimizing (over the available data) the “distance” between the approximate probability distribution and the true one; in probability theory, the distance between two distributions is the so-called Kullback-Leibler (KL) divergence [6]. Various families of approximate distributions can be used, e.g., mixture models [2].
We have ready access to tools to draw samples from simple distributions: e.g., the randn function in many programming languages generates samples of a one-dimensional Gaussian distribution with desired parameters, and we can create a histogram of the samples to verify this. However, for more exotic distributions in higher-dimensional spaces, we need an algorithm. A naive approach would be a grid search: we systematically evaluate the probability at every sample point on an $n$-dimensional grid of points. However, this would waste most of its effort on areas with low probability, and this wastefulness gets worse as $n$ increases—the so-called curse of dimensionality. Instead, we can rely on Markov chain Monte Carlo (MCMC), a class of methods that construct a chain of samples that concentrate in the high-probability regions and approximate a desired distribution. (We also typically only keep every $L$th chain entry to ensure the samples are independent.) A feature of MCMC that makes it particularly helpful is that it only compares (i.e., considers the ratio of) the probability of a candidate sample point with that of a previous one to decide whether to accept it in the chain. Since the probability distribution we care about comes from Bayes’ theorem, this means we can ignore the common scaling factor and simply use (7): we only need to evaluate the prior and the likelihood, and the latter only requires that we evaluate the observation model $h(\mathsf{x})$ of each candidate sample point. The Metropolis-Hastings method [7] is the basic workhorse method in this class. If we also have access to the gradient of the observation operator, then we can use a hybrid (or Hamiltonian) Monte Carlo [2] to accelerate the sampling process. And importantly, if we suspect that there are multiple high-probability regions (i.e., multiple solutions of our inverse problem), then we can use the Method of Parallel Tempering [8], which uses multiple chains to more thoroughly explore the state space to identify these regions.
@@ 2.3.2 Example: static estimation of point vortices from pressure
For a fluid dynamics demonstration of Bayesian inference with a nonlinear observation operator, we will consider a simple but illustrative example: using a small number of pressure sensors to estimate the instantaneous position and strength of one or more point vortices in an otherwise quiescent unbounded two-dimensional fluid domain. This example is covered in more detail by Eldredge and Le Provost [9]. The state vector $\mathsf{x}$ comprises $n=3N$ entries: the $(x_{\mathsf{J}},y_{\mathsf{J}})$ coordinates and strength $\Gamma_{\mathsf{J}}$ of each of the $N$ point vortices. (We will assume we know $N$.) The problem is nondimensionalized by the fluid density and reference strength $\Gamma_{\text{ref}}$ and length $L_{\text{ref}}$. The observation vector $\mathsf{y}$ consists of the $d$ pressure measurements (relative to ambient pressure), $\Delta p_{\alpha}$, $\alpha=1,\ldots,d$, at $d$ sensor locations. The observation operator $h$ in this example is obtained analytically from the Bernoulli equation; it depends nonlinearly on all state components, quadratically on the strengths and inverse algebraically on the positions. Importantly, a pair of vortices, if comparably close to each other as they are to the sensors, generates a coupled component of the pressure that depends on the product of their strengths. The true measurement vector $\mathsf{y}^{\ast}$ in this example is synthetically generated from the observation operator applied to the true state, $\mathsf{x}^{\ast}$, and adding Gaussian noise.
@@
Figure 3: MCMC samples in the position plane (left) and expected vorticity (right) of a single point vortex estimated from three pressure sensors (brown squares). The black dot and gray ellipse depicts, respectively, the true state and the covariance evaluated there. Adapted with permission from [9].
@@
Figure 3 depicts the MCMC samples in the position plane and the resulting expectation of vorticity using a true observation from three noisy pressure sensors (with noise standard deviation of $5\times 10^{-4}$). The estimator is successful in estimating the location and strength of the vortex, though it has some uncertainty about the vertical position of the vortex. This uncertainty can also be visualized with the posterior covariance ellipse shown with the samples. This covariance was obtained analytically in this case by computing (10) with $H\approx\nabla h(\mathsf{x}^{\ast})$, where $\mathsf{x}^{\ast}$ is the true state, and infinite prior covariance. Computing this covariance analytically obviously assumes that we know the true state, but it is useful at least for illuminating the information available in the observation operator and its gradient in a nonlinear problem. For example, it can be used to show that the maximum uncertainty of estimating a vortex grows quickly as the distance between the vortex and the sensors increases [9].
@@
Figure 4: MCMC samples in the position plane of a single point vortex estimated from two pressure sensors (brown squares). Circle is the manifold of possible states, and black dot is the true state. Adapted with permission from [9].
@@
In the previous case, we had enough sensors to infer the vortex’s characteristics. With only two sensors, we likely lack sufficient data to uniquely determine the vortex. Indeed, there is a manifold of equally-plausible states, and it can be shown that this manifold takes the form of a helix whose axis is parallel to the $\Gamma$ axis. In Figure 4, the MCMC samples are shown in the vortex position plane to be distributed about the projection of this manifold, a circle. We have no further knowledge to estimate the expected state. Another way to interpret this behavior is through $H\approx\nabla h(\mathsf{x}^{\ast})$. Similar to the manometer example, this matrix is rank deficient, in this case along a direction tangent to the manifold at $\mathsf{x}^{\ast}$. If the vortex state is perturbed along this manifold, the two sensors will not detect any change. In the manometer example, we simply fell back on the prior to estimate the uncertainty along the rank-deficient direction. Unfortunately, for the vortex, the prior covariance is infinite, so the posterior uncertainty remains infinite along the rank-deficient direction. We need more information about the vortex (via the prior) or another sensor.
@@
With the number of sensors at least as large as the number of states, $d\geq n$, we generally avoid issues of non-uniqueness in estimating a single point vortex. Suppose the vortex is not a point vortex, but has a spatially-distributed vortex core. Can we also estimate the vortex radius if we simply add a fourth pressure sensor? Unfortunately, estimating the size of a vortex from pressure data is problematic for two reasons [9]. First, the effect of a vortex’s size on its pressure distribution is only appreciable if the sensors lie within the radius of the vortex core: the uncertainty grows like the inverse cube of the radius when the sensors are outside of this core. Second, even when the sensors lie inside the core, they have trouble distinguishing the states from one another, so the uncertainty remains large. Note that, as long as we do not seek to estimate the vortex size, the first behavior is actually useful: we can reliably estimate the position and strength of a real vortex with an observation model based on an idealized point vortex, as long as the real vortex is not very large.
@@
Figure 5: Expected vorticity of two point vortices estimated from 8 sensors (left) and three point vortices from 11 sensors (right). Adapted with permission from [9].
@@ 2.4 Comments on static estimation
Spectral decomposition of the observation operator is important
@@
In estimation problems with nonlinear observation models, we would like to exploit the same benefits of the spectral decomposition of $H$ to break down the estimation process. The obvious candidate for this decomposition is the Jacobian $\nabla h$, based on a Taylor expansion of the observation model. In probability theory, the relevant tool is the Fisher information matrix, which provides a general measure of the amount of information about the state $\mathsf{x}$ that is available in the observation $\mathsf{y}$. For a Gaussian distribution of $\mathsf{y}$, the Fisher information is given by $\nabla h(\mathsf{x})^{T}\Sigma_{\mathcal{E}}^{-1}\nabla h(\mathsf{x})$. But at which state do we evaluate this matrix? In the vortex example, we evaluated it at the true state, but this is hardly a useful strategy since we generally do not know this state. Cui and Zahm [10] propose that we take the expected value of this Fisher information over the prior, resulting in the state-space Gramian,
@@
In the case of a linear observation model, decomposed as in (14), this Gramian reduces to $C_{\mathsf{x}}=VS^{T}SV^{T}$. In other words, the spectral decomposition of $H$ automatically provides us with the eigendecomposition of $C_{\mathsf{x}}$, whose eigenvalues are in the diagonal matrix $\Lambda_{\mathsf{x}}=S^{T}S$ (and clearly non-negative) and whose eigenvectors are the columns of $V$. Spantini et al. [11] showed that the eigenvectors with the largest eigenvalues indicate the state directions in which measurements are most informative relative to the prior—that is, the dominant directions in which the state will be updated by an observation. This suggests that, for a nonlinear observation operator, we can compute the eigenvectors $V$ and eigenvalues $\Lambda_{\mathsf{x}}$ of $C_{\mathsf{x}}$ defined by (24) and reap the same benefits as for a linear operator. If we only have access to an ensemble of prior samples of the state, then (24) can be approximated by a simple average over the ensemble, as shown in (63).
What about the basis $U$ of the measurement space? Le Provost et al. [12] suggest an analogous measurement-space Gramian,
@@
A good prior is very useful
@@ 3 Sequential estimation of flows
This distribution over $\mathsf{x}_{1}$ then serves as a prior for folding in a new observation $\mathsf{y}_{1}$ with Bayes theorem (6), as we did before, to obtain a posterior distribution,
@@
At the end of this two-step process, we have a distribution of states conditioned on all previous measurements, and clearly this process can be repeated indefinitely. These equations constitute the recursive nonlinear filtering equations. In the context of filtering, the first step is called the prediction, or forecast; the second is called the update, or analysis.
In the discussion that follows, we will generally assume that the forecast and observation operators are derived from physics. However, they can also be learned from data, and we will provide a more thorough discussion of estimation and uncertainty with learned operators in Section 4. However, the data can also be used to learn the transition probabilities and likelihood directly, and we highlight here an aerodynamics example of this approach in the recent cluster-based Bayesian model by Kaiser et al. [13]. In that work, the goal was to estimate the force on a delta wing from the measurement of a few surface pressure sensors. They assembled a large data set from experimental measurements of surface pressure and force on a delta wing at different angles of attack and undergoing various maneuvers. The force and pressure measurements form a joint state space, in which they first identified clusters, and from the data sequences, inferred the transition probability from each cluster to another. With these state transitions, and with a Gaussian model for the likelihood of a pressure observation belonging to a certain cluster, they used Bayes theorem to estimate the posterior distribution of force at each time step.
@@ 3.1 Linear operators, Gaussian noise
For linear forecast $F_{k}$ and observation $H_{k}$ models and Gaussian process and measurement noise, $\eta\sim\mathcal{N}(0,\Sigma_{\eta})$ and $\varepsilon\sim\mathcal{N}(0,\Sigma_{\mathcal{E}})$, the state distributions themselves remain Gaussian, and the filtering equations reduce exactly to equations for the state mean and covariance for both the forecast and the analysis, just as they did for the Bayesian inference problem on its own. This is the classical Kalman filter [14], with a forecast step
@@
An example of a use of a linear Kalman filter for unsteady aerodynamics estimation is provided by An et al. [15]. In that work, the objective was to use sparse surface pressure measurements to estimate the lift coefficient on a NACA 0009 airfoil in a steady free stream undergoing large-amplitude pitch oscillations. The state vector comprised the lift coefficient and the pressure sensor values; including the measurements in the state ensured that the observation matrix simply picked off these measurement entries to predict the observation. A linear forecast operator for the lift coefficient was constructed from a modified Goman–Khrabrov model [16] for the lift, combined with a least-squares regression for the lift from the pressure measurements, trained from a single case.

### v3

@@ 2.3.1 Sampling the posterior
So how do we proceed? If our goal is just to find the expected state given the true measurement, then we should note that this is the $\mathsf{x}$ that minimizes (22). This note effectively frames the state estimation as an optimization problem, called the maximum a posteriori (MAP) estimate, and we could use optimization tools to solve it. However, our goal is also to characterize the posterior uncertainty, so we still seek the posterior distribution in some form. We have a few options. If our prior distribution is already reasonably precise, then we do not have to move very far from it to obtain the posterior. In such a case we can linearize the observation operator about the prior mean and then use the Kalman update that we presented in the previous section, taking $H\approx\nabla h(\mu_{\mathsf{x}})$. We will revisit this idea later when we discuss sequential estimation. However, if our prior has a lot of uncertainty, or we suspect that the solution $\mathsf{x}$ to (1) is not unique, then we need a more thorough strategy, and we can get this by sampling the posterior. Sampling from a distribution provides us with an indirect tool for characterizing it: with an ensemble $\{\mathsf{x}^{i}\}_{i=1}^{M}$ of $M$ independent samples, each drawn from the posterior distribution $\mathsf{x}^{i}\sim\pi(\mathsf{x}|\mathsf{y}^{\ast})$, we can approximately compute expected values by summing over the samples,
@@
We can also use the samples to develop approximate probabilistic models of the posterior, $\tilde{\pi}(\mathsf{x}|\mathsf{y})$, to do further work, e.g., to serve as a generative model for drawing new instances of the state. There are many approaches for carrying out this approximation task, most of which consist of minimizing (over the available data) the “distance” between the approximate probability distribution and the true one; in probability theory, the distance between two distributions is the so-called Kullback-Leibler (KL) divergence [kullback1951information]. Various families of approximate distributions can be used, e.g., mixture models [bishop2006pattern].
We have ready access to tools to draw samples from simple distributions: e.g., the randn function in many programming languages generates samples of a one-dimensional Gaussian distribution with desired parameters, and we can create a histogram of the samples to verify this. However, for more exotic distributions in higher-dimensional spaces, we need an algorithm. A naive approach would be a grid search: we systematically evaluate the probability at every sample point on an $n$-dimensional grid of points. However, this would waste most of its effort on areas with low probability, and this wastefulness gets worse as $n$ increases—the so-called curse of dimensionality. Instead, we can rely on Markov chain Monte Carlo (MCMC), a class of methods that construct a chain of samples that concentrate in the high-probability regions and approximate a desired distribution. (We also typically only keep every $L$th chain entry to ensure the samples are independent.) A feature of MCMC that makes it particularly helpful is that it only compares (i.e., considers the ratio of) the probability of a candidate sample point with that of a previous one to decide whether to accept it in the chain. Since the probability distribution we care about comes from Bayes’ theorem, this means we can ignore the common scaling factor and simply use (7): we only need to evaluate the prior and the likelihood, and the latter only requires that we evaluate the observation model $h(\mathsf{x})$ of each candidate sample point. The Metropolis-Hastings method [Chib1995] is the basic workhorse method in this class and is very simple to implement. If we also have access to the gradient of the observation operator, then we can use a hybrid (or Hamiltonian) Monte Carlo [bishop2006pattern] to accelerate the sampling process. And importantly, if we suspect that there are multiple high-probability regions (i.e., multiple solutions of our inverse problem), then we can use the Method of Parallel Tempering [Sambridge2014], which uses multiple chains to more thoroughly explore the state space to identify these regions.
@@ 2.3.2 Example: static estimation of point vortices from pressure
For a fluid dynamics demonstration of Bayesian inference with a nonlinear observation operator, we will consider a simple but illustrative example: using a small number of pressure sensors to estimate the instantaneous position and strength of one or more point vortices in an otherwise quiescent unbounded two-dimensional fluid domain. This example is covered in more detail by Eldredge and Le Provost [eldredge2024bayesian]. The state vector $\mathsf{x}$ comprises $n=3N$ entries: the $(x_{\mathsf{J}},y_{\mathsf{J}})$ coordinates and strength $\Gamma_{\mathsf{J}}$ of each of the NN point vortices. (We will assume we know $N$.) The problem is nondimensionalized by the fluid density and reference strength $\Gamma_{\text{ref}}$ and length $L_{\text{ref}}$. The observation vector $\mathsf{y}$ consists of the $d$ pressure measurements (relative to ambient pressure), $\Delta p_{\alpha}$, $\alpha=1,\ldots,d$, at $d$ sensor locations. The observation operator $h$ in this example is obtained analytically from the Bernoulli equation; it depends nonlinearly on all state components, quadratically on the strengths and inverse algebraically on the positions. Importantly, a pair of vortices, if comparably close to each other as they are to the sensors, generates a coupled component of the pressure that depends on the product of their strengths. The true measurement vector $\mathsf{y}^{\ast}$ in this example is synthetically generated from the observation operator applied to the true state, $\mathsf{x}^{\ast}$, and adding Gaussian noise.
@@
Figure 3: One point vortex estimated from three pressure sensors (brown squares). (Left) MCMC samples (light blue) in position plane, with true state (black dot) and covariance ellipse; (right) expected vorticity. Adapted with permission from Eldredge and Le Provost (J. Fluid Mech., 2024) [eldredge2024bayesian], copyright 2024 Jeff D. Eldredge.
@@
Figure 3 depicts the MCMC samples in the position plane and the resulting expectation of vorticity using a true observation from three noisy pressure sensors (with noise standard deviation of $5\times 10^{-4}$). The estimator is successful in estimating the location and strength of the vortex, though it has some uncertainty about the vertical position of the vortex. This uncertainty can also be visualized with the posterior covariance ellipse shown with the samples. This covariance was obtained analytically in this case by computing (10) with $H\approx\nabla h(\mathsf{x}^{\ast})$, where $\mathsf{x}^{\ast}$ is the true state, and infinite prior covariance. Computing this covariance analytically obviously assumes that we know the true state, but it is useful at least for illuminating the information available in the observation operator and its gradient in a nonlinear problem. For example, it can be used to show that the maximum uncertainty of estimating a vortex grows quickly as the distance between the vortex and the sensors increases [eldredge2024bayesian].
@@
Figure 4: MCMC samples in the position plane of a single point vortex estimated from two pressure sensors (brown squares). Circle is the manifold of possible states, and black dot is the true state. Adapted with permission from Eldredge and Le Provost (J. Fluid Mech., 2024) [eldredge2024bayesian], copyright 2024 Jeff D. Eldredge.
@@
In the previous case, we had enough sensors to infer the vortex’s characteristics. With only two sensors, we likely lack sufficient data to uniquely determine the vortex. Indeed, there is a manifold of equally-plausible states, and it can be shown that this manifold takes the form of a helix whose axis is parallel to the $\Gamma$ axis. In Figure 4, the MCMC samples are shown in the vortex position plane to be distributed about the projection of this manifold, a circle. We have no further knowledge to estimate the expected state. Another way to interpret this behavior is through $H\approx\nabla h(\mathsf{x}^{\ast})$. Similar to the manometer example, this matrix is rank deficient, in this case along a direction tangent to the manifold at $\mathsf{x}^{\ast}$. If the vortex state is perturbed along this manifold, the two sensors will not detect any change. In the manometer example, we simply fell back on the prior to estimate the uncertainty along the rank-deficient direction. Unfortunately, for the vortex, the prior covariance is infinitely large, so the posterior uncertainty remains infinite along the rank-deficient direction. We need more information about the vortex (via a better prior) or another sensor.
@@
With the number of sensors at least as large as the number of states, $d\geq n$, we generally avoid issues of non-uniqueness in estimating a single point vortex. Suppose the vortex is not a point vortex, but has a spatially-distributed vortex core. Can we also estimate the vortex radius if we simply add a fourth pressure sensor? Unfortunately, estimating the size of a vortex from pressure data is problematic for two reasons [eldredge2024bayesian]. First, the effect of a vortex’s size on its pressure distribution is only appreciable if the sensors lie within the radius of the vortex core: the uncertainty grows like the inverse cube of the radius when the sensors are outside of this core. Second, even when the sensors lie inside the core, they have trouble distinguishing the states from one another, so the uncertainty remains large. Note that, as long as we do not seek to estimate the vortex size, the first behavior is actually useful: we can reliably estimate the position and strength of a real vortex with an observation model based on an idealized point vortex, as long as the real vortex is not very large in extent.
@@
Figure 5: Expected vorticity of two point vortices estimated from 8 sensors (left) and three point vortices from $1$ sensors (right). Adapted with permission from Eldredge and Le Provost (J. Fluid Mech., 2024) [eldredge2024bayesian], copyright 2024 Jeff D. Eldredge.
@@ 2.4 Comments on static estimation
Spectral decomposition of the observation operator is important.
@@
In estimation problems with nonlinear observation models, we would like to exploit the same benefits of the spectral decomposition of $H$ to break down the estimation process. The obvious candidate for this decomposition is the Jacobian $\nabla h$, based on a Taylor expansion of the observation model. In probability theory, the relevant tool is the Fisher information matrix, which provides a general measure of the amount of information about the state $\mathsf{x}$ that is available in the observation $\mathsf{y}$. For a Gaussian distribution of $\mathsf{y}$, the Fisher information is given by $\nabla h(\mathsf{x})^{T}\Sigma_{\mathcal{E}}^{-1}\nabla h(\mathsf{x})$. But at which state do we evaluate this matrix? In the vortex example, we evaluated it at the true state, but this is hardly a useful strategy since we generally do not know this state. Cui and Zahm [cui2021data] propose that we take the expected value of this Fisher information over the prior, resulting in the state-space Gramian,
@@
In the case of a linear observation model, decomposed as in (14), this Gramian reduces to $C_{\mathsf{x}}=VS^{T}SV^{T}$. In other words, the spectral decomposition of $H$ automatically provides us with the eigendecomposition of $C_{\mathsf{x}}$, whose eigenvalues are in the diagonal matrix $\Lambda_{\mathsf{x}}=S^{T}S$ (and clearly non-negative) and whose eigenvectors are the columns of $V$. Spantini et al. [spantini2015optimal] showed that the eigenvectors with the largest eigenvalues indicate the state directions in which measurements are most informative relative to the prior—that is, the dominant directions in which the state will be updated by an observation. This suggests that, for a nonlinear observation operator, we can compute the eigenvectors $V$ and eigenvalues $\Lambda_{\mathsf{x}}$ of $C_{\mathsf{x}}$ defined by (24) and reap the same benefits as for a linear operator. If we only have access to an ensemble of prior samples of the state, then (24) can be approximated by a simple average over the ensemble, as shown in (63).
What about the basis $U$ of the measurement space? Le Provost et al. [le2022low] suggest an analogous measurement-space Gramian,
@@
A good prior is very useful.
@@ 3 Sequential estimation of flows
This distribution over $\mathsf{x}_{1}$ then serves as a prior for folding in a new observation $\mathsf{y}_{1}$ with Bayes’ theorem (6), as we did before, to obtain a posterior distribution,
@@
At the end of this two-step process, we have a distribution of states conditioned on all previous measurements, and clearly this process can be repeated indefinitely. These equations constitute an exact general framework for advancing the distribution, usually called recursive Bayesian estimation (or a Bayes filter) [arulampalam2002tutorial]. In the context of filtering, the first step is called the prediction, or forecast; the second is called the update, or analysis.
In the discussion that follows, we will generally assume that the forecast and observation operators are derived from physics. However, they can also be learned from data, and we will provide a more thorough discussion of estimation and uncertainty with learned operators in Section 4. However, the data can also be used to learn the transition probabilities and likelihood directly, and we highlight here an aerodynamics example of this approach in the recent cluster-based Bayesian model by Kaiser et al. [kaiser2024cluster]. In that work, the goal was to estimate the force on a delta wing from the measurement of a few surface pressure sensors. They assembled a large data set from experimental measurements of surface pressure and force on a delta wing at different angles of attack and undergoing various maneuvers. The force and pressure measurements form a joint state space, in which they first identified clusters, and from the data sequences, inferred the transition probability from each cluster to another. With these state transitions, and with a Gaussian model for the likelihood of a pressure observation belonging to a certain cluster, they used Bayes’ theorem to estimate the posterior distribution of force at each time step.
@@ 3.1 Linear operators, Gaussian noise
For linear models of forecast $F_{k}$ and observation $H_{k}$ and Gaussian process and measurement noise, $\eta\sim\mathcal{N}(0,\Sigma_{\eta})$ and $\varepsilon\sim\mathcal{N}(0,\Sigma_{\mathcal{E}})$, the state distributions themselves remain Gaussian, and the filtering equations reduce exactly to equations for the state mean and covariance for both the forecast and the analysis, just as they did for the Bayesian inference problem on its own. This is the classical Kalman filter [kalman1960new], with a forecast step
@@
An example of a use of a linear Kalman filter for unsteady aerodynamics estimation is provided by An et al. [an2021lift]. In that work, the objective was to use sparse surface pressure measurements to estimate the lift coefficient on a NACA 0009 airfoil in a steady free stream undergoing large-amplitude pitch oscillations. The state vector comprised the lift coefficient and the pressure sensor values; including the measurements in the state ensured that the observation matrix simply picked off these measurement entries to predict the observation. A linear forecast operator for the lift coefficient was constructed from a modified Goman–Khrabrov model [goman1994state] for the lift, combined with a least-squares regression for the lift from the pressure measurements, trained from a single case.